# Harnessing Textual Semantic Priors for Knowledge Transfer and Refinement in CLIP-Driven Continual Learning

### v1

@@ Abstract
Continual learning (CL) aims to equip models with the ability to learn from a stream of tasks without forgetting previous knowledge. With the progress of vision-language models like Contrastive Language-Image Pre-training (CLIP), their promise for CL has attracted increasing attention due to their strong generalizability. However, the potential of rich textual semantic priors in CLIP in addressing the stability–plasticity dilemma remains underexplored. During backbone training, most approaches transfer past knowledge without considering semantic relevance, leading to interference from unrelated tasks that disrupt the balance between stability and plasticity. Besides, while text-based classifiers provide strong generalization, they suffer from limited plasticity due to the inherent modality gap in CLIP. Visual classifiers help bridge this gap, but their prototypes lack rich and precise semantics. To address these challenges, we propose Semantic-Enriched Continual Adaptation (SECA), a unified framework that harnesses the anti-forgetting and structured nature of textual priors to guide semantic-aware knowledge transfer in the backbone and reinforce the semantic structure of the visual classifier. Specifically, a Semantic-Guided Adaptive Knowledge Transfer (SG-AKT) module is proposed to assess new images’ relevance to diverse historical visual knowledge via textual cues, and aggregate relevant knowledge in an instance-adaptive manner as distillation signals. Moreover, a Semantic-Enhanced Visual Prototype Refinement (SE-VPR) module is introduced to refine visual prototypes using inter-class semantic relations captured in class-wise textual embeddings. Extensive experiments on multiple benchmarks validate the effectiveness of our approach. Code is available in the supplementary materials.
@@ Introduction
Continual learning (CL) aims to equip models with the ability to learn continuously from a stream of tasks (Lu et al. 2024), a key requirement for adapting to evolving open-world scenarios. A fundamental challenge is the stability-plasticity dilemma (Kim and Han 2023; Chen et al. 2023), which requires models to preserve previously learned knowledge (stability) while adapting to new information (plasticity). We investigate the challenging rehearsal-free Class-Incremental Learning (CIL) setting, where the model must continually learn new classes without rehearsal and predict over all seen categories without task identities.
Recent advances in vision-language models, such as Contrastive Language-Image Pre-training (CLIP) (Radford et al. 2021), provide strong zero-shot capabilities and serve as promising backbones for CL. While several methods (Zhou et al. 2025; Huang et al. 2024; Jha, Gong, and Yao 2024) leverage CLIP mainly as a powerful backbone, the potential of its textual semantic priors for improving the stability–plasticity trade-off remains underexplored. To retain old knowledge in backbones, most approaches (Kirkpatrick et al. 2017; Kang, Park, and Han 2022; Asadi et al. 2023) rely on regularization or distillation to enforce consistency with model after the most recent task. However, such non-selective knowledge retention can lead to semantic interference, where unrelated or noisy knowledge from previous tasks hampers adaptation to new tasks and distorts earlier feature spaces. As shown in Fig. 1(a), when learning a new class like ’cat’, prior knowledge from similar classes such as ’dog’ or ’bird’ can be helpful, while unrelated classes like ’vehicle’ introduce interference that degrades learning. As for classifiers, although the text-only designs (Huang et al. 2024; Jha, Gong, and Yao 2024) offer strong generalization, they limit plasticity due to the inherent modality gap in CLIP. Prior studies (Liang et al. 2022; Fahim, Murphy, and Fyshe 2024) attributes this gap to a geometric separation in the shared embedding space, where visual and textual embeddings remain distinctly apart in the feature space. This modality gap is further exacerbated in CL, as task-isolated training prevents joint calibration, reducing the text-based classifier’s ability to adapt to new classes.
@@
Continual Learning. Continual Learning (CL) aims to incrementally acquire new knowledge while preserving performance on previously learned tasks. Traditional CL methods typically learn models from scratch and can be categorized into three types: (a) replay-based methods (Bonicelli et al. 2022; Aljundi et al. 2019a, b), which store exemplar samples to jointly train with new tasks; (b) regularization-based methods (Jung et al. 2020; Zenke, Poole, and Ganguli 2017; Aljundi et al. 2018; Wang et al. 2025), which constrain parameter updates through additional regularization terms; and (c) expansion-based methods (Hung et al. 2019; Li et al. 2019), which dynamically allocate new parameters for future tasks to improve model plasticity. With the advent of large-scale pre-trained models (Dosovitskiy 2020; Radford et al. 2021), recent studies (Wang et al. 2022b; Lu et al. 2024; Liang and Li 2024; He et al. 2025) have explored Parameter-Efficient Fine-Tuning (PEFT) as a promising solution for CL, enabling task adaptation without full model retraining. In particular, CLIP-based continual learning (Huang et al. 2024; Zhou et al. 2025; Jha, Gong, and Yao 2024; Menabue et al. 2024) has attracted increasing interest due to its strong zero-shot generalization. STAR (Menabue et al. 2024) treats the pre-trained CLIP as a powerful task-id selector. RAPF (Huang et al. 2024) introduces a lightweight learnable projector and a selective pseudo-feature replay approach, enhances fine-grained class discrimination. PROOF (Zhou et al. 2025) improves image-text alignment via a learnable cross-attention module and alleviates forgetting through a weight averaging strategy.
Parameter-Efficient Fine-Tuning (PEFT) for Vision-Lanugage Models. Recent vision-language models (VLMs) such as CLIP (Radford et al. 2021) and BLIP (Li et al. 2022) enable strong zero-shot transfer but suffer from domain shift on downstream tasks. To address this, PEFT methods adapt VLMs by updating only a small set of parameters. Prompt tuning-based methods ($e.g.$, CoOp (Lester, Al-Rfou, and Constant 2021), CoCoOp (Zhou et al. 2022), ProDA (Lu et al. 2022), KgCoOp (Yao, Zhang, and Xu 2023)) learn soft prompts to condition the text branch, while adapter-based approaches (e.g., CLIP-Adapter (Gao et al. 2024), Tip-Adapter (Zhang et al. 2021)) introduce lightweight projection matrices or cache-based classifiers to improve model performance on downstream tasks. Other PEFT strategies include low-rank tuning (LoRA (Hu et al. 2021)), prefix tuning (Li and Liang 2021), and Mixture-of-Experts (MoE (Jacobs et al. 1991)). Despite tuning only a small subset of parameters, these methods have demonstrated effectiveness comparable to full-model fine-tuning.
@@
where $[\mathbf{P}^{s}]_{m}$ is the $m$-th learnable token in $\mathbf{P}^{s}$. This sequence is then fed into the frozen text encoder $\mathcal{F}_{\mathrm{T}}(\cdot)$ to generate the textual feature for $y$: $\mathcal{F}_{\mathrm{T}}(y;\mathbf{P}^{s})=\mathcal{F}_{\mathrm{T}}(\mathbf{E}_{y}^{s})\in\mathbb{R}^{d_{\mathrm{T}}}$.
CLIP Fine-Tuning with Visual Adapters. To adapt the visual encoder $\mathcal{F}_{\mathrm{V}}(\cdot)$ to downstream tasks, we introduce a set of task-shared lightweight adapters (Chen et al. 2022) into the visual backbone. At the $s$-th task, we denote the adapter set as $\mathcal{A}=\{\mathcal{A}_{l}\}_{l=1}^{L}$, where each $\mathcal{A}_{l}$ is inserted alongside the $l$-th transformer block. Given an input $\hat{\mathbf{h}}_{l}$ of the Feed-Forward Network (FFN) at layer $l$, the output is computed as:
@@
[…]
@@
where $\phi(\cdot)$ denotes the LayerNorm (Ba, Kiros, and Hinton 2016) operation to to stabilize training. $\alpha_{x}^{(p)}$ represents the relevance of its ground-truth class $y$ to learned visual knowledge from $p$-th adapter in shared semantic space. The relevance scores then used to guide the aggregation of visual knowledge $\mathbf{V}^{(x)}$ from the adapter pool:
@@
[…]
@@
Table 1: Experimental results on ImageNetR and ImageNetA. We report the averaged results over 3 trials. VPT-NSP and RAPF denote the feature replay-free version of VPT-NSP (Lu et al. 2024) and RAPF (Huang et al. 2024), respectively. The highest results are in bold, and the second best results are underlined.
Method Venue Backbone ER/FR 10S-ImageNetR 10S-ImageNetA 20S-ImageNetR 20S-ImageNetA Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ L2P (Wang et al. 2022b) CVPR-22 ViT-21k 72.34 77.36 44.04 51.24 69.64 75.28 40.48 49.62 CODA (Smith et al. 2023) CVPR-23 ViT-21k 73.31 78.47 52.08 63.92 69.96 75.34 44.62 54.86 RanPAC (McDonnell et al. 2023) NeurIPS-23 ViT-21k 77.90 82.91 62.40 67.58 Adam-adapter (Zhou et al. 2024) IJCV-24 ViT-21k 65.79 72.42 48.81 58.84 57.42 64.75 48.65 59.55 SSIAT (Tan et al. 2024) CVPR-24 ViT-21k 79.38 83.63 62.43 70.83 75.67 82.30 59.16 68.45 VPT-NSP (Lu et al. 2024) NeurIPS-24 ViT-21k 77.95 83.44 53.83 63.93 75.69 81.87 49.81 61.41 DIA (Li et al. 2024) CVPR-25 ViT-21k 79.03 85.61 59.78 70.43 76.32 83.51 ACMap (Fukuda, Kera, and Kawamoto 2024) CVPR-25 ViT-21k 73.50 79.50 56.19 65.19 ZS-CLIP CLIP 74.93 81.56 47.33 58.35 74.93 82.09 47.33 59.36 L2P (Wang et al. 2022b) CVPR-22 CLIP 75.98 81.67 47.86 59.35 68.78 76.87 47.54 59.77 DualPrompt (Wang et al. 2022a) CVPR-22 CLIP 75.77 82.01 48.18 59.05 69.41 77.07 48.05 60.22 CODA (Smith et al. 2023) CVPR-22 CLIP 67.52 78.00 50.24 64.32 64.53 75.23 49.95 65.08 Adam-adapter (Zhou et al. 2024) IJCV-24 CLIP 71.35 78.65 59.35 68.56 68.75 76.71 58.55 67.72 RAPF (Huang et al. 2024) ECCV-24 CLIP 73.23 82.20 45.54 60.67 71.28 81.66 43.85 58.54 RAPF (Huang et al. 2024) ECCV-24 CLIP 79.62 86.28 55.37 67.32 80.28 85.58 49.85 65.28 CLAP (Jha, Gong, and Yao 2024) NeurIPS-24 CLIP 79.98 85.77 58.66 69.35 79.12 85.03 55.84 67.72 VPT-NSP (Lu et al. 2024) NeurIPS-24 CLIP 77.45 83.60 47.14 59.04 77.52 83.94 47.07 59.88 VPT-NSP (Lu et al. 2024) NeurIPS-24 CLIP 82.48 87.94 61.42 71.76 82.06 88.09 60.70 72.57 PROOF (Zhou et al. 2025) T-PAMI-25 CLIP 77.25 82.69 55.67 65.50 77.05 82.83 54.05 64.53 SECA (Ours) CLIP 83.18 88.58 65.09 74.45 82.25 88.19 62.80 73.02 SECA++ (Ours) CLIP 83.41 88.75 65.77 74.65 83.02 88.62 64.41 74.64
@@
Method ER/FR 10S-CIFAR100 20S-CIFAR100 Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ ZS-CLIP 66.68 75.15 66.68 75.93 L2P (Wang et al. 2022b) 73.08 81.90 68.67 79.18 DualPrompt (Wang et al. 2022a) 72.51 81.45 69.91 79.74 CODA (Smith et al. 2023) 62.25 76.98 41.98 69.78 SLCA (Zhang et al. 2023) 67.58 80.53 66.84 78.96 Adam-adapter (Zhou et al. 2024) 65.50 75.76 58.12 70.18 RAPF (Huang et al. 2024) 79.04 86.19 79.26 86.87 CLAP (Jha, Gong, and Yao 2024) 78.21 86.13 77.35 86.08 PROOF (Zhou et al. 2025) 76.29 84.88 76.13 85.12 SECA (Ours) 79.79 86.70 77.73 85.35 SECA++ (Ours) 81.59 87.80 80.07 87.11
Datasets. We evaluate our method on three representative CIL benchmarks: ImageNetR (Hendrycks et al. 2021a), ImageNetA (Hendrycks et al. 2021b), and CIFAR-100 (Krizhevsky, Hinton et al. 2009). ImageNetR comprises 30,000 images spanning 200 classes, covering a wide range of visual styles. ImageNetA contains 7,500 challenging real-world adversarial examples images across 200 classes. CIFAR100 is a commonly-used dataset in continual learning, which consists of 60000 32$\times$32 images of 100 classes. We adopt both 10-split and 20-split settings, where the class set is evenly divided into 10 or 20 non-overlapping tasks.
Evaluation Metrics. Following prior PEFT-based CIL works (Tan et al. 2024; Gao, Cen, and Chang 2024; Huang et al. 2024), we adopt two standard evaluation metrics: (1) The last session accuracy (Last): the final accuracy over all classes after completing the last task, and (2) Average accuracy (Avg), the average accuracy across all incremental tasks. All experiments are conducted using three random seeds, and we report the mean performances.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. We adopt CLIP (Radford et al. 2021) with a ViT-B/16 (Dosovitskiy 2020) visual encoder pre-trained by OPENAI as our backbone. Our model is optimized using Adam (Kingma and Ba 2014) with an initial learning rate of 0.001. For the training schedule, each task is trained for 10 epochs on ImageNetR and ImageNetA, and 3 epochs on CIFAR100. The batch size is set to 64 for ImageNetR and ImageNetA and 100 for CIFAR100. The capacity of the frozen adapter pool $\mathcal{P}$ in SG-AKT is set to 5, and the temperature factor $\tau^{\prime}$ in $\mathcal{L}_{\mathrm{SG-AKT}}$ is set to 0.05.
SECA++: An Enhanced Variant of SECA. Following existing PEFT-based CIL methods (Tan et al. 2024; Huang et al. 2024; Lu et al. 2024), we further implement an enhanced variant, denoted as SECA++, which incorporates Gaussian sampling-based feature replay to alleviate inter-task confusion during the training of multi-modal classifier, particularly for semantically confusing or fine-grained classes.
@@
We compare SECA with recent state-of-the-art PEFT-based approaches, including prompt-based methods (Wang et al. 2022b; Lu et al. 2024; Smith et al. 2023), adapter-based methods (Li et al. 2024; Tan et al. 2024; Huang et al. 2024), the zero-shot CLIP (ZS-CLIP) and other representative techniques built upon either ViT-B/16 pretrained on ImageNet-21K (Russakovsky et al. 2015) (ViT-21K) or CLIP-pretrained backbones. Evaluations are conducted on ImageNetR, ImageNetA and CIFAR100, and the results are shown in Tab.LABEL:tab:comparison_imagenet and Tab.2. ‘ER/FP’ indicates the use of experience replay (Rolnick et al. 2019) or distributional sampling-based feature replay. Notably, most CLIP-based CIL methods incorporate such replay strategies into their frameworks. For a fair comparison, we reproduce the results of two representative methods, RAPF (Huang et al. 2024) and VPT-NSP (Lu et al. 2024), using a CLIP backbone under a replay-free setting. These variants are denoted as RAPF and VPT-NSP.
@@
Idx H-PEFT SG-AKT SE-VPR 10S-ImageNetA 10S-CIFAR100 10S-ImageNetR Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ 1 - - - 47.33 58.35 67.19 76.23 74.93 81.56 4 - - 55.78$\pm$1.36 67.97$\pm$0.38 73.97$\pm$0.65 82.75$\pm$0.28 80.57$\pm$0.15 87.04$\pm$0.35 5 - 57.91$\pm$0.55 68.56$\pm$0.66 75.93$\pm$0.09 84.16$\pm$0.21 81.36$\pm$0.49 87.55$\pm$0.29 6 - 62.62$\pm$0.65 73.15$\pm$0.25 77.13$\pm$0.56 84.91$\pm$0.43 81.30$\pm$0.55 87.63$\pm$0.12 7 65.09$\pm$0.48 74.45$\pm$0.76 79.79$\pm$0.42 86.70$\pm$0.16 83.18$\pm$0.34 88.58$\pm$0.07
@@
Method 10S-ImageNetA 10S-CIFAR100 Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Methods without SE-VPR. Seq. 55.78$\pm$1.36 67.97$\pm$0.38 73.97$\pm$0.65 82.75$\pm$0.28 CLIP-KD 55.01$\pm$0.68 67.00$\pm$0.27 71.21$\pm$0.61 80.98$\pm$0.50 Vanilla 57.05$\pm$0.75 67.93$\pm$0.47 74.55$\pm$0.26 83.26$\pm$0.32 Avg-KD 56.90$\pm$0.11 68.02$\pm$0.76 75.05$\pm$0.78 83.65$\pm$0.33 SG-AKT (Ours) 57.91$\pm$0.55 68.56$\pm$0.66 75.93$\pm$0.09 84.16$\pm$0.21 Methods with SE-VPR. Seq. 62.62$\pm$0.65 73.15$\pm$0.25 77.13$\pm$0.56 84.91$\pm$0.43 CLIP-KD 62.39$\pm$0.77 73.01$\pm$1.17 73.30$\pm$0.80 83.76$\pm$0.18 Vanilla 63.51$\pm$0.71 73.89$\pm$0.69 76.42$\pm$0.44 85.75$\pm$0.37 Avg-KD 64.32$\pm$0.59 74.08$\pm$0.82 78.98$\pm$0.51 86.47$\pm$0.30 SG-AKT (Ours) 65.09$\pm$0.48 74.45$\pm$0.76 79.79$\pm$0.42 86.70$\pm$0.16
SG-AKT v.s. other Distillation Strategies. We compare our SG-AKT with four alternative training strategies in Tab.4: (1) Seq. denotes sequential tuning without any distillation loss; (2) CLIP-KD distills from the frozen CLIP model as a global teacher; (3) Vanilla distills knowledge from the most recent task model, following prior works (Kang, Park, and Han 2022; Asadi et al. 2023); (4) Avg-KD is a simplified variant of SG-AKT that aggregates adapter pool features by averaging, without text-aware attention. Tab. 4 shows that SG-AKT consistently outperforms all variants, with +0.77% and +1.58% Last-acc improvements over Avg-KD and Vanilla on 10S-ImageNetA, respectively. Although these two designs are effective for knowledge retention, they fall short compared to SG-AKT, highlighting the effectiveness of leveraging textual information to prioritize semantically relevant transfer.

### v3

@@ Abstract
Continual learning (CL) aims to equip models with the ability to learn from a stream of tasks without forgetting previous knowledge. With the progress of vision-language models like Contrastive Language-Image Pre-training (CLIP), their promise for CL has attracted increasing attention due to their strong generalizability. However, the potential of rich textual semantic priors in CLIP in addressing the stability–plasticity dilemma remains underexplored. During backbone training, most approaches transfer past knowledge without considering semantic relevance, leading to interference from unrelated tasks that disrupt the balance between stability and plasticity. Besides, while text-based classifiers provide strong generalization, they suffer from limited plasticity due to the inherent modality gap in CLIP. Visual classifiers help bridge this gap, but their prototypes lack rich and precise semantics. To address these challenges, we propose Semantic-Enriched Continual Adaptation (SECA), a unified framework that harnesses the anti-forgetting and structured nature of textual priors to guide semantic-aware knowledge transfer in the backbone and reinforce the semantic structure of the visual classifier. Specifically, a Semantic-Guided Adaptive Knowledge Transfer (SG-AKT) module is proposed to assess new images’ relevance to diverse historical visual knowledge via textual cues, and aggregate relevant knowledge in an instance-adaptive manner as distillation signals. Moreover, a Semantic-Enhanced Visual Prototype Refinement (SE-VPR) module is introduced to refine visual prototypes using inter-class semantic relations captured in class-wise textual embeddings. Extensive experiments on multiple benchmarks validate the effectiveness of our approach. Our code is available at: https://github.com/HHHLF/SECA_master.
@@ Introduction
Continual learning (CL) aims to equip models with the ability to learn continuously from a stream of tasks (vptnspNIPS-24), a key requirement for adapting to evolving open-world scenarios. A fundamental challenge is the stability-plasticity dilemma (kim2023stability; chen2023stability), which requires models to preserve previously learned knowledge (stability) while adapting to new information (plasticity). We investigate the challenging rehearsal-free Class-Incremental Learning (CIL) setting, where the model must continually learn new classes without rehearsal and predict over all seen categories without task identities.
Recent advances in vision-language models, such as Contrastive Language-Image Pre-training (CLIP) (2021clip), provide strong zero-shot capabilities and serve as promising backbones for CL. While several methods (PROOF-TPAMI25; rapfECCV24; clap4clip-NIPS24) leverage CLIP mainly as a powerful backbone, the potential of its textual semantic priors for improving the stability–plasticity trade-off remains underexplored. To retain old knowledge in backbones, most approaches (EwC; AFC-CVPR22; PSRD-ICML23) rely on regularization or distillation to enforce consistency with model after the most recent task. However, such non-selective knowledge retention can lead to semantic interference, where unrelated or noisy knowledge from previous tasks hampers adaptation to new tasks and distorts earlier feature spaces. As shown in Fig. 1(a), when learning a new class like ’cat’, prior knowledge from similar classes such as ’dog’ or ’bird’ can be helpful, while unrelated classes like ’vehicle’ introduce interference that degrades learning. As for classifiers, although the text-only designs (rapfECCV24; clap4clip-NIPS24) offer strong generalization, they limit plasticity due to the inherent modality gap in CLIP. Prior studies (liang2022mind; fahim2024s-its-not-a) attributes this gap to a geometric separation in the shared embedding space, where visual and textual embeddings remain distinctly apart in the feature space. This modality gap is further exacerbated in CL, as task-isolated training prevents joint calibration, reducing the text-based classifier’s ability to adapt to new classes.
@@
Continual Learning. Continual Learning (CL) aims to incrementally acquire new knowledge while preserving performance on previously learned tasks. Traditional CL methods typically learn models from scratch and can be categorized into three types: (a) replay-based methods (replay1; replay2; replay3), which store exemplar samples to jointly train with new tasks; (b) regularization-based methods (regularization1; regularization2; regularization3; wanghuaijie2025ekpc), which constrain parameter updates through additional regularization terms; and (c) expansion-based methods (expansion1; expansion2), which dynamically allocate new parameters for future tasks to improve model plasticity. With the advent of large-scale pre-trained models (ViT; 2021clip), recent studies (l2pCVPR22; vptnspNIPS-24; infloraCVPR24; helingfeng2025ckaa) have explored Parameter-Efficient Fine-Tuning (PEFT) as a promising solution for CL, enabling task adaptation without full model retraining. In particular, CLIP-based continual learning (rapfECCV24; PROOF-TPAMI25; clap4clip-NIPS24; menabue2024semantic) has attracted increasing interest due to its strong zero-shot generalization. STAR (menabue2024semantic) treats the pre-trained CLIP as a powerful task-id selector. RAPF (rapfECCV24) introduces a lightweight learnable projector and a selective pseudo-feature replay approach, enhances fine-grained class discrimination. PROOF (PROOF-TPAMI25) improves image-text alignment via a learnable cross-attention module and alleviates forgetting through a weight averaging strategy.
Parameter-Efficient Fine-Tuning (PEFT) for Vision-Lanugage Models. Recent vision-language models (VLMs) such as CLIP (2021clip) and BLIP (BLIP) enable strong zero-shot transfer but suffer from domain shift on downstream tasks. To address this, PEFT methods adapt VLMs by updating only a small set of parameters. Prompt tuning-based methods ($e.g.$, CoOp (PromptTuning), CoCoOp (cocoop), ProDA (proda), KgCoOp (KG-CoOp)) learn soft prompts to condition the text branch, while adapter-based approaches (e.g., CLIP-Adapter (CLIP-adapter), Tip-Adapter (TIP-adapter)) introduce lightweight projection matrices or cache-based classifiers to improve model performance on downstream tasks. Other PEFT strategies include low-rank tuning (LoRA (lora)), prefix tuning (PrefixTuning), and Mixture-of-Experts (MoE (Mixture-of-Experts)). Despite tuning only a small subset of parameters, these methods have demonstrated effectiveness comparable to full-model fine-tuning.
@@
where $\mathbf{P}^{s}]_{m}$ is the $m$-th learnable token in $\mathbf{P}^{s}$. This sequence is then fed into the frozen text encoder $\mathcal{F}_{\mathrm{T}}(\cdot)$ to generate the textual feature for $y$: $\mathcal{F}_{\mathrm{T}}(y;\mathbf{P}^{s})=\mathcal{F}_{\mathrm{T}}(\mathbf{E}_{y}^{s})\in\mathbb{R}^{d_{\mathrm{T}}}$.
CLIP Fine-Tuning with Visual Adapters. To adapt the visual encoder $\mathcal{F}_{\mathrm{V}}(\cdot)$ to downstream tasks, we introduce a set of task-shared lightweight adapters (adapter) into the visual backbone. At the $s$-th task, we denote the adapter set as $\mathcal{A}=\{\mathcal{A}_{l}\}_{l=1}^{L}$, where each $\mathcal{A}_{l}$ is inserted alongside the $l$-th transformer block. Given an input $\hat{\mathbf{h}}_{l}$ of the Feed-Forward Network (FFN) at layer $l$, the output is computed as:
@@
where $\tau$ is a temperature factor and $\langle\cdot,\cdot\rangle$ denotes the cosine similarity between $\ell_{2}$-normalized visual and textual features. Given a data point $(\mathbf{x},y)\in\mathcal{D}^{s}$, the learnable PEFT modules are trained through a standard cross-entropy loss:
@@
where $\phi(\cdot)$ denotes the LayerNorm (Layernorm) operation to to stabilize training. $\alpha_{x}^{(p)}$ represents the relevance of its ground-truth class $y$ to learned visual knowledge from $p$-th adapter in shared semantic space. The relevance scores then used to guide the aggregation of visual knowledge $\mathbf{V}^{(x)}$ from the adapter pool:
@@
where $\hat{\mathbf{c}}_{\mathrm{V},k}^{s-1}$ is the refined visual prototype after the (s1)(s-1)-th task of the previous class $k$. By injecting textual semantics into visual prototypes and preserving temporal consistency, SE-VPR establishes a powerful visual-side classifier to compensate for the modality gap.
@@
Table 1: Experimental results on ImageNetR and ImageNetA. We report the averaged results over 3 trials. VPT-NSP and RAPF denote the feature replay-free version of VPT-NSP (vptnspNIPS-24) and RAPF (rapfECCV24), respectively. The highest results are in bold, and the second best results are underlined.
Method Venue Backbone ER/FR 10S-ImageNetR 10S-ImageNetA 20S-ImageNetR 20S-ImageNetA Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ L2P (l2pCVPR22) CVPR-22 ViT-21k 72.34 77.36 44.04 51.24 69.64 75.28 40.48 49.62 CODA (codaCVPR23) CVPR-23 ViT-21k 73.31 78.47 52.08 63.92 69.96 75.34 44.62 54.86 RanPAC (ranpac-NIPS23) NeurIPS-23 ViT-21k 77.90 82.91 62.40 67.58 Adam-adapter (adamIJCV2024) IJCV-24 ViT-21k 65.79 72.42 48.81 58.84 57.42 64.75 48.65 59.55 SSIAT (ssaitCVPR24) CVPR-24 ViT-21k 79.38 83.63 62.43 70.83 75.67 82.30 59.16 68.45 VPT-NSP (vptnspNIPS-24) NeurIPS-24 ViT-21k 77.95 83.44 53.83 63.93 75.69 81.87 49.81 61.41 DIA (DIA-CVPR2025) CVPR-25 ViT-21k 79.03 85.61 59.78 70.43 76.32 83.51 ACMap (ACMap-CVPR2025) CVPR-25 ViT-21k 73.50 79.50 56.19 65.19 ZS-CLIP CLIP 74.93 81.56 47.33 58.35 74.93 82.09 47.33 59.36 L2P (l2pCVPR22) CVPR-22 CLIP 75.98 81.67 47.86 59.35 68.78 76.87 47.54 59.77 DualPrompt (dualpromptECCV22) CVPR-22 CLIP 75.77 82.01 48.18 59.05 69.41 77.07 48.05 60.22 CODA (codaCVPR23) CVPR-22 CLIP 67.52 78.00 50.24 64.32 64.53 75.23 49.95 65.08 Adam-adapter (adamIJCV2024) IJCV-24 CLIP 71.35 78.65 59.35 68.56 68.75 76.71 58.55 67.72 RAPF (rapfECCV24) ECCV-24 CLIP 73.23 82.20 45.54 60.67 71.28 81.66 43.85 58.54 RAPF (rapfECCV24) ECCV-24 CLIP 79.62 86.28 55.37 67.32 80.28 85.58 49.85 65.28 CLAP (clap4clip-NIPS24) NeurIPS-24 CLIP 79.98 85.77 58.66 69.35 79.12 85.03 55.84 67.72 VPT-NSP (vptnspNIPS-24) NeurIPS-24 CLIP 77.45 83.60 47.14 59.04 77.52 83.94 47.07 59.88 VPT-NSP (vptnspNIPS-24) NeurIPS-24 CLIP 82.48 87.94 61.42 71.76 82.06 88.09 60.70 72.57 PROOF (PROOF-TPAMI25) T-PAMI-25 CLIP 77.25 82.69 55.67 65.50 77.05 82.83 54.05 64.53 SECA (Ours) CLIP 83.18 88.58 65.09 74.45 82.25 88.19 62.80 73.02 SECA++ (Ours) CLIP 83.41 88.75 65.77 74.65 83.02 88.62 64.41 74.64
@@
Method ER/FR 10S-CIFAR100 20S-CIFAR100 Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ ZS-CLIP 66.68 75.15 66.68 75.93 L2P (l2pCVPR22) 73.08 81.90 68.67 79.18 DualPrompt (dualpromptECCV22) 72.51 81.45 69.91 79.74 CODA (codaCVPR23) 62.25 76.98 41.98 69.78 SLCA (slcaICCV23) 67.58 80.53 66.84 78.96 Adam-adapter (adamIJCV2024) 65.50 75.76 58.12 70.18 RAPF (rapfECCV24) 79.04 86.19 79.26 86.87 CLAP (clap4clip-NIPS24) 78.21 86.13 77.35 86.08 PROOF (PROOF-TPAMI25) 76.29 84.88 76.13 85.12 SECA (Ours) 79.79 86.70 77.73 85.35 SECA++ (Ours) 81.59 87.80 80.07 87.11
Datasets. We evaluate our method on three representative CIL benchmarks: ImageNetR (imagenet-r), ImageNetA (imagenet-a), and CIFAR-100 (cifar100). ImageNetR comprises 30,000 images spanning 200 classes, covering a wide range of visual styles. ImageNetA contains 7,500 challenging real-world adversarial examples images across 200 classes. CIFAR100 is a commonly-used dataset in continual learning, which consists of 60000 $\times32$ images of 100 classes. We adopt both 10-split and 20-split settings, where the class set is evenly divided into 10 or 20 non-overlapping tasks.
Evaluation Metrics. Following prior PEFT-based CIL works (ssaitCVPR24; cpromptCVPR24; rapfECCV24), we adopt two standard evaluation metrics: (1) The last session accuracy (Last): the final accuracy over all classes after completing the last task, and (2) Average accuracy (Avg), the average accuracy across all incremental tasks. All experiments are conducted using three random seeds, and we report the mean performances.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. We adopt CLIP (2021clip) with a ViT-B/16 (ViT) visual encoder pre-trained by OPENAI as our backbone. Our model is optimized using Adam (adam_optimizer) with an initial learning rate of 0.001. For the training schedule, each task is trained for 10 epochs on ImageNetR and ImageNetA, and 3 epochs on CIFAR100. The batch size is set to 64 for ImageNetR and ImageNetA and 100 for CIFAR100. The capacity of the frozen adapter pool $\mathcal{P}$ in SG-AKT is set to 5, and the temperature factor $\tau^{\prime}$ in $\mathcal{L}_{\mathrm{SG-AKT}}$ is set to 0.05.
SECA++: An Enhanced Variant of SECA. Following existing PEFT-based CIL methods (ssaitCVPR24; rapfECCV24; vptnspNIPS-24), we further implement an enhanced variant, denoted as SECA++, which incorporates Gaussian sampling-based feature replay to alleviate inter-task confusion during the training of multi-modal classifier, particularly for semantically confusing or fine-grained classes.
@@
We compare SECA with recent state-of-the-art PEFT-based approaches, including prompt-based methods (l2pCVPR22; vptnspNIPS-24; codaCVPR23), adapter-based methods (DIA-CVPR2025; ssaitCVPR24; rapfECCV24), the zero-shot CLIP (ZS-CLIP) and other representative techniques built upon either ViT-B/16 pretrained on ImageNet-21K (imagenet-21k) (ViT-21K) or CLIP-pretrained backbones. Evaluations are conducted on ImageNetR, ImageNetA and CIFAR100, and the results are shown in Tab.LABEL:tab:comparison_imagenet and Tab.2. ‘ER/FP’ indicates the use of experience replay (ER) or distributional sampling-based feature replay. Notably, most CLIP-based CIL methods incorporate such replay strategies into their frameworks. For a fair comparison, we reproduce the results of two representative methods, RAPF (rapfECCV24) and VPT-NSP (vptnspNIPS-24), using a CLIP backbone under a replay-free setting. These variants are denoted as RAPF and VPT-NSP.
@@
[…]
@@
[…]
SG-AKT v.s. other Distillation Strategies. We compare our SG-AKT with four alternative training strategies in Tab.4: (1) Seq. denotes sequential tuning without any distillation loss; (2) CLIP-KD distills from the frozen CLIP model as a global teacher; (3) Vanilla distills knowledge from the most recent task model, following prior works (AFC-CVPR22; PSRD-ICML23); (4) Avg-KD is a simplified variant of SG-AKT that aggregates adapter pool features by averaging, without text-aware attention. Tab. 4 shows that SG-AKT consistently outperforms all variants, with +0.77% and +1.58% Last-acc improvements over Avg-KD and Vanilla on 10S-ImageNetA, respectively. Although these two designs are effective for knowledge retention, they fall short compared to SG-AKT, highlighting the effectiveness of leveraging textual information to prioritize semantically relevant transfer.